OpenStudy (anonymous):

Write the equation of the line that passes by (–1, 4) with a slope of –1 in standard form.

OpenStudy (amistre64):

standard form is: ax+by=c ; where slope is defined as -a/b

OpenStudy (amistre64):

x + y = c gives us a slope of -1/1 = -1 use the point to calibrate for c

OpenStudy (anonymous):

So It Would Be A(-1)+B(4)=3 ?

OpenStudy (anonymous):

x+y=3

OpenStudy (amistre64):

x + y = c (-1) + (4) = c 3 = c x+y = 3

OpenStudy (anonymous):

So What's The Standform Of The Equation ?

OpenStudy (anonymous):

Standard Form*

OpenStudy (anonymous):

you want to set up the equation like this. (y-y1)=m(x-x1) where x1,y1 are your given coordinates. and m is they slope. so you just plug and chug

OpenStudy (amistre64):

i would assume it is the equation which we gave you and worked out for you ...

OpenStudy (anonymous):

y-(4) =-1(x-(-1)) y-4=-x-1 y=-x+3 that is the line of the equation

OpenStudy (anonymous):

Well Its Not . . Nevermind Someone Else Gave Me The RIGHT Answer

OpenStudy (anonymous):

hope that helped...?

OpenStudy (anonymous):

It Did Thanks :D

OpenStudy (amistre64):

http://www.mathwords.com/s/standard_form_for_the_equation_of_a_line.htm then you should complain to whoever wrote your question :)

OpenStudy (anonymous):

The standard equation of line is Ax+By=C, here A=B=1,C=3 therefore 1x+1y=3 is the standard equation

OpenStudy (amistre64):

slope intercept form of a line is: y = -x + 3